Please answer this question for me.
A local bakery determines that the probability of a customer ordering a blueberry muffin is 50%. 5 customers come into the bakery.
A. What is the probability that 3 out of the 5 order a blueberry muffin?
B. Explain why this scenario would meet the criteria of a binomial distribution.


Answer by stanbon(75887)

Please answer this question for me.
A local bakery determines that the probability of a customer ordering a blueberry muffin is 50%. 5 customers come into the bakery.
A. What is the probability that 3 out of the 5 order a blueberry muffin?
P(x=3) = 5C3(0.5)^3(0.5)^2 = 10/32 = 5/16
---------------------
B. Explain why this scenario would meet the criteria of a binomial distribution.
n=5 independent trials
p= 1/2 on each trial
===============
Cheers,
Stan H.